OpenStudy (anonymous):

Inez wants to have $15,000 in 2 years. Use the present value formula to calculate how much Inez should invest now at 6 % interest, compounded quarterly in order to reach her goal. a. $13,349.95 b. $13,315.65 c. $ 12,231.90 d. $ 11,877.98

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Can't open that at all

OpenStudy (anonymous):

It was only ur quiz. :( Couldn't find any thing for you srry

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Can I have a medal for my attempt?

OpenStudy (anonymous):

Ty!!!

OpenStudy (anonymous):

Future = (Present) (1.06)^2 Plug and chug.

OpenStudy (ranga):

Here is the compound interest formula: \[\Large A = P(1 + \frac{ r }{ n })^{nt}\]A = Amount at maturity = $15,000 P = Principal Amount = ? r = Annual interest rate in decimal = 6% = 0.06 n = compounding period (compounded how many times a year) = 4 (quarterly compounding) t = years invested = 2 Solve for P

OpenStudy (anonymous):

How would I solve this

OpenStudy (anonymous):

15000 = x (1.06)^2

OpenStudy (ranga):

15000 = P( 1 + 0.06/4)^(4*2) = P(1.015)^8 P = 15000 / (1.015)^8 = ?

OpenStudy (anonymous):

$13,315.65

OpenStudy (ranga):

@douglaswinslowcooper Does your method take into account quarterly compounding?

OpenStudy (anonymous):

No, I missed that detail entirely. My apologies. You were right and I was wrong.

OpenStudy (ranga):

np @douglaswinslowcooper @tashasp Yes, $13,315.67